Here is the problem.. "A motorcycle traveling at 50 miles per hour overtakes a car traveling at 30 miles per hour that had a three-hour head start. How far from the starting point are the two vehicles?"
Answer by mananth(16946)

motorcycle speed = 50 mph
car = 30 mph
travelling same direction
so effective speed = 50-30=20
head start 3 hours
distance = 3*30 =90 miles
t=d/r
catch up time = 90/20
=4.5 hours
speed of motorcycle = 50 mph.
time = 4.5 hours
so distance = 50*4.5 = 225 miles. from start
m.ananth@hotmail.ca